Fiber technology Inc., manufactures glass fibers used in the communications industry, The company's materials and parts manager is currently revising the inventory policy for XL-20, one of the chemicals used in the production process. The chemical is purchased in 10 pound canisters for $110 each. The firm uses 5,400 canisters per year. The controller estimates that it costs $165 to place and recieve a typical order of XL-20. The annual cost of storing XL-20 is $5.50 per canister. 2. Use EOQ formula to determine the optimal order quantity 3. What is the total annual cost of ordering and storing XL-20 at the economic order quantity? 4. How many orders will be placed per year? 5. Fiber Technology’s controller, Jay Turnbull, recently attended a seminar on JIT purchasing. Afterward he analyzed the cost of storing XL-20, including the costs of wasted space and inefficiency. He was shocked when he concluded that the real annual holding cost was $34.20 per canister. Turnbull then met with Doug Kaplan, Fiber Technology’s purchasing manager. Together they contacted Reno Industries, the supplier of XL-20, about a JIT purchasing arrangement. After some discussion and negotiation, Kaplan concluded that the cost of placing an order for XL-20 could be reduced to just $35.00. Using these new cost estimates, Turnbull computed the new EOQ for XL-20. .. use the equation approach to compute the new EOQ …how many orders will be placed per year?
Answer #1

Answer 2.

EOQ = square root of (2 × annual demand × Cost Per Order / Carrying Cost Per unit)

EOQ = square root of ((2 × 5400 × 165) / 5.50)

EOQ=569 units

Answer 4.

Number of orders to be placed per year = 5400 / 569 = 9.49

Answer 3.

total annual cost of ordering = 9*165

   1485

total annual cost of storing = [(0+569)/2]*5.50

   1565

total annual cost of ordering and storing

3050

Answer 5.

EOQ = square root of (2 × annual demand × Cost Per Order / Carrying Cost Per unit)

EOQ = square root of ((2 × 5400 × 35) / 34.20)

EOQ= 105 units

Number of orders to be placed per year = 5400 / 105 = 51 orders